Compute the divergence and curl of the vector point functions. 𝐹 = 𝑋 ^2π‘Œπ‘π‘– βˆ’ 2𝑋 𝑍^ 3 𝑗 + 𝑋 𝑍^ 2π‘˜.

1
Expert's answer
2022-03-20T06:43:45-0400

"divF=\\frac{\\partial F_1}{\\partial x}+\\frac{\\partial F_2}{\\partial y}+\\frac{\\partial F_3}{\\partial z}=2xyz-0+2xz=2xz\\left( y+1 \\right) \\\\rotF=\\left| \\begin{matrix}\ti&\t\tj&\t\tk\\\\\t\\frac{\\partial}{\\partial x}&\t\t\\frac{\\partial}{\\partial y}&\t\t\\frac{\\partial}{\\partial z}\\\\\tx^2yz&\t\t-2xz^3&\t\txz^2\\\\\\end{matrix} \\right|=i\\left( 0+6xz^2 \\right) -j\\left( z^2-x^2y \\right) +k\\left( -2z^3-x^2z \\right) =\\\\=\\left( 6xz^2,x^2y-z^2,-x^2z-2z^3 \\right)"